# GmNet: Revisiting Gating Mechanisms From A Frequency View

### v2

@@ 7 Conclusion
In this paper, we systematically explore the effect of gating mechanisms on the training dynamics of neural networks from a frequency perspective. We provide both theoretical and experimental analysis to reveal how the element-wise product and activation interplay to manage the balance between different frequencies learning. The element-wise product can broaden the spectrum of the signals while introduce high-frequency noise. The activation function can effectively control the frequency output based on its own characters. Benefited by the comprehensive study on GLUs, we introduce a lightweight model named GmNet to efficiently utilize the information of various frequency. It shows impressive performance on both effectiveness and efficiency.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We clearly state our contributions and scope in the abstract and introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitations in Sec. 6
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We have provided the full set of assumptions and a complete (and correct) proof.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We have included all details of our implementation both in main text and appendix.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [N/A]
  24. Justification: The data used in this paper are all public and we will release our code upon acceptance
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We have included details of our implementation in Sec. 5
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We cover the details in Sec. 5 about evaluation setups.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We have provided details of our compute resources in the Sec. 5.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Yes, we follow all the code of ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: They are in Sec. 6
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our work has no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We have cited the original papers of the datasets.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: We do not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Our paper does not involve crowdsourcing or research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our paper does not involve crowdsourcing or research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: We do not use LLM for help.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v3

@@ 6 Conclusion
This paper tackled the prevalent low-frequency bias in lightweight networks through a novel frequency-based analysis of gating mechanisms. We found that in a Gated Linear Unit (GLU), element-wise multiplication introduces valuable high-frequency information, while the paired activation function provides crucial control to filter for useful signals over noise. Our resulting model, the Gating Mechanism Network (GmNet), validates this approach by setting a new state-of-the-art in efficient network design. This work demonstrates that a frequency-aware methodology is a promising path toward creating future models that are both efficient and representationally robust.
7 Statements
7.1 Ethics Statement
In our paper, we strictly follow the ICLR ethical research standards and laws. To the best of our knowledge, our work abides by the General Ethical Principles.
7.2 Reproducibility Statement
We adhere to ICLR reproducibility standards and ensure the reproducibility of our work. All datasets we employed are publicly available. We will provide the code to reviewers and area chairs in the supplementary material.
Appendix A Appendix
A.1 Implementation Details
A.1.1 Pseudo-codes of Model Architectures
In our modified ResNet18, featured in Fig. 2, we adjust the activation function as the different variants. As an example, we provide the pseudo-codes of Res18-Gate-ReLU in the Algorithm 1
Algorithm 1 Pseudo-codes of Res18-Gate-ReLU
def Block(x, in_planes, planes)
  out = Conv2d(x, in_planes, planes, 3, 1, 1)
  out = BatchNorm2d(x)
  out = ReLU(out) * out
  out = Conv2d(out, planes, planes, 3, 1, 1)
  out = BatchNorm2d(out)
  out += self.shortcut(x)
  out = ReLU(out)
  return out
For the proposed GmNet, featured in Fig. 4, we provide the pseudo-code of GmNet in the Algorithm 2. Also, for ease of reproduction, we include a separate file in the supplementary materials dedicated to our GmNet.
Algorithm 2 Pseudo-codes of GmNet
def Block(x, dim, mlp_ratio)
  input = x
  x = DWConv2d(x, dim, dim, 7, 1, 3, group=dim)
  x = BatchNorm2d(x)
  x = Conv2d(x, dim, mlp_ratio*dim, 1)
  x = ReLU6(x) * x
  x = Conv2d(x, mlp_ratio*dim, dim, 1)
  x = BatchNorm2d(x)
  x = DWConv2d(x, dim, dim, 7, 1, 3, group=dim)
  x = input + drop_path(layer_scale(x))
  return x
A.1.2 Training Recipes
We first provide the detailed training settings of variants of ResNet18 in Table 7.
A.1.3 GmNet Variants
We provide the setting of variants of GmNet in Table. 6.
Variant $C_{1}$ depth ratio Params FLOPs
GmNet-S1 40 $[2,2,10,2]$ $[3,3,3,2]$ 3.7 M 0.6 G
GmNet-S2 48 $[2,2,8,3]$ $[3,3,3,2]$ 6.2 M 0.9 G
GmNet-S3 48 $[3,3,8,3]$ $[4,4,4,4]$ 7.8 M 1.2 G
GmNet-S4 68 $[3,3,11,3]$ $[4,4,4,4]$ 17.0 M 2.7 G
config value
image size 32
optimizer SGD
base learning rate $0.1$
weight decay 5e-4
optimizer momentum $0.9$
batch size 128
learning rate schedule cosine decay
training epochs 100
Table 7: Res18 variants training setting
We also provide a detailed training configures of GmNet in this section as shown in the Table 8.
config value
image size 224
optimizer AdamW
base learning rate $3\mathrm{e}-3$
weight decay 0.03
optimizer momentum $\beta_{1},\beta_{2}=0.9,0.999$
batch size 2048
learning rate schedule cosine decay
warmup epochs 5
training epochs 300
AutoAugment rand-m1-mstd0.5-inc1
label smoothing 0.1
cutmix 0.4
color jitter 0.
drop path $0.(\mathrm{S}1/\mathrm{S}2),0.02(\mathrm{S}3/\mathrm{~S}4)$
Table 8: GmNet training setting
Figure 8: Top-1 accuracy vs Latency on A100. Our models have significantly smaller latency compared to related works.
Table 9: Comparison of different GLU designs for GmNet-S3 on ImageNet-1K.
GLUs Top-1 Params GPU $r=12$ $r=24$ $r=36$ $r=48$ $r=60$
(%) (M) (ms) Low High Low High Low High Low High Low High
$\sigma(x)\cdot\text{FC}(x)$ 79.2 20.2 3.6 10.8 51.4 39.6 8.7 52.6 4.4 62.9 3.4 69.9 2.4
$\sigma(\text{FC}(x))\cdot x$ 79.6 20.2 3.4 9.4 48.9 35.0 9.1 51.1 3.6 62.1 3.1 68.7 2.4
A.2 More GLUs designs.
In this section, we present additional block designs to demonstrate the efficiency and effectiveness of our proposed architecture. As shown in Table 9, we conducted experiments incorporating fully connected (FC) layers into the Gated Linear Units (GLUs). Adding an extra FC layer to one branch of the GLUs may slightly improve performance. However, this modification significantly increases the number of parameters and reduces the model’s latency.
Moreover, this enhanced architecture does not perform better in classifying different frequency components. The reason is that features processed by an FC layer cannot effectively emphasize various frequency components. While adding more parameters and employing different training methods might enhance the capability to learn different frequency components, in lightweight models, the simplest GLU design often delivers better performance. This observation is consistent with findings from many recent studies on lightweight models.
A.3 More comparisons of the main results.
We provide more comparisons of the main results in Fig. 8. We plot a larger latency-acc trade-off figure to include more methods including EdgeViT, MobileNetV2 and GhostNet.
A.4 Results on CUB-100.
We evaluated GmNet-S1 on the CUB-100 dataset. The results are competitive, as shown in Table 10. Compared to ShuffleNet-V2, GmNet-S1 achieves better performance with a smaller model size.
Variant Params (M) Top-1 Acc (%)
ShuffleNet-V2 3.5 76.0
GmNet-S1 3.1 81.5
Table 10: Results on CUB-100.
A.5 Performance with SwiGLU and SiLU
We have conducted experiments with the SiLU variant on CIFAR-10 to illustrate the learning dynamics in Fig. 14. Following the reviewer’s advice, we adapted both SwiGLU and SiLU to GmNet-S3 and trained the models under the same settings. The results in Table 11 show that both activations improve performance compared to the baseline configuration, which removes activations inside the GLU. However, their gains remain consistently smaller than those achieved by our proposed design, confirming that the proposed GmNet block is better aligned with the frequency characteristics and architectural constraints of lightweight models.
Model / Activation Top-1 Acc (%)
Baseline (Identity) 70.5
+ SiLU 77.9
+ SwiGLU 77.2
+ Proposed (ReLU6-GLU) 79.3
Table 11: Results of using SiLU and SwiGLU in GmNet-S3.
A.6 Performance of changing ReLU6 to ReLU.
Table 12 shows the performance change when replacing ReLU6 with ReLU across multiple GmNet variants. The degradation is consistent across all model sizes.
Variant ReLU6 $\rightarrow$ ReLU
GmNet-S1 75.5 $\rightarrow$ 74.7
GmNet-S2 78.3 $\rightarrow$ 77.4
GmNet-S3 79.3 $\rightarrow$ 78.3
GmNet-S4 81.5 $\rightarrow$ 80.5
Table 12: Performance degradation when replacing ReLU6 with ReLU.
A.7 Studies on aliasing and robustness
We train Res18-Gate-ReLU and Res18-Gate-GELU on CIFAR-10 following the setting of ref2 to show how different activation functions affect robustness. We conducted an adversarial robustness ablation under PGD attacks. As shown below, the ReLU-based GLU—which emphasizes higher-frequency components—exhibits slightly lower PGD accuracy:
  • Res18-Gate-ReLU: 45.7% PGD accuracy
  • Res18-Gate-GELU: 46.7% PGD accuracy
Clean accuracy remains similar for both variants (82.2% vs. 82.8%), but the $\sim$1% drop in PGD robustness for the ReLU gate is consistent with prior findings that stronger high-frequency reliance increases vulnerability to aliasing and adversarial perturbations.
A.8 Experiments on ConvNeXt
To show the generalization of gating structure, we adapted our GLU to ConvNeXt-Tiny and trained the model using the same setting as GmNet on ImageNet-1k. As shown in Table 13, GLU improves the performance without introducing any additional computational cost.
The improvement is smaller compared to GmNet, which may result from the fact that large CNNs are less affected by activation-induced frequency bias. Block design should be architecture-dependent.
Model Top-1 Acc (%) Params (M) FLOPs (G)
ConvNeXt 82.5 28.6 4.46
ConvNeXt + GLU 83.0 28.6 4.46
Table 13: Comparison between ConvNeXt-Tiny and its GLU-enhanced variant.
A.9 More results on Efformerv2 and MobileNetV2.
Figure 9: The illustration of modifications of MobileNetV2 and EfficientFormer-V2.
We first show the illustration of the modifications for both models in . Moreover, we plot the testing curves with different settings of thresholds. As shown in Fig. 13, the overall performance trend is consistent with the charts and figures presented in the main text. The model using ReLU6 shows better performance on the high-frequency components where the overall performance also surpasses other models. Meanwhile, the model using GELU performs better on low-frequencies. The extra results demonstrate the effect of GLU on helping model learning various frequency information and the improvements on high-frequency information have more impact on the final performance.
Figure 10: Additional results under different threshold configurations of different variants of ResNet18.
Figure 11: Additional results under different threshold configurations of different variants of MobileNetv2.
Figure 12: Additional results under different threshold configurations of different variants of Efficientformer-V2.
Figure 13: Additional results under different threshold configurations of different variants of MobileNetv2 and EfficientFormer-v2 (Eformerv2).
A.10 Quantitative spectral evidences.
To further explore the effect of different activation functions, we provide the quantitative spectral evidence. Based on the GmNet-S3 variants, we computed high/low-frequency energy ratios across multiple layers and model variants as shown in Table 14. Firstly, we extract the layers before/after the gate and compute their high/low-frequency energy ratios to show the spectral changes. Also, to demonstrate how different activation functions affect the model’s frequency response, we compute the high/low-frequency energy ratios of the first DW-Conv layers in each stages.
Table 14: High/Low Frequency Ratio Comparison. We computed the spectral changes from the layer before to the layer after the gate which are defined as f and g respectively. We also compute the high/low-frequency energy ratios of the first 7×7 DW-Conv layers of each stage.
High/Low Frequency Ratio Changes before/after the gate variants
Stage Layer Pair ReLU6 (H/L) GELU (H/L) ReLU (H/L)
0.1 f $\rightarrow$ g 0.1195 $\rightarrow$ 0.1200 0.0575 $\rightarrow$ 0.0040 0.5172 $\rightarrow$ 0.5674
1.1 f $\rightarrow$ g 0.0989 $\rightarrow$ 0.1429 0.0423 $\rightarrow$ 0.0022 0.1422 $\rightarrow$ 0.1751
2.1 f $\rightarrow$ g 0.0386 $\rightarrow$ 0.0706 0.0252 $\rightarrow$ 0.0013 0.0018 $\rightarrow$ 0.0030
3.1 f $\rightarrow$ g 0.0019 $\rightarrow$ 0.0281 0.0106 $\rightarrow$ 0.0003 0.0032 $\rightarrow$ 0.0341
High/Low Frequency Ratio Comparison at DW-Conv layers
Stage Layer GELU (H/L) ReLU (H/L) ReLU6 (H/L)
0.1 1st DW-Conv 0.1203 1.1553 0.7057
1.1 1st DW-Conv 0.0695 0.3040 0.2452
2.1 1st DW-Conv 0.0381 0.0024 0.0697
3.1 1st DW-Conv 0.0057 0.0088 0.0142
We define the low frequencies as the central 1/4 region of the 2D spectrum. Table 14 shows that a consistent spectral pattern distinguishing smooth and non-smooth activations. For GELU, the transition from f to g typically increases the low-frequency response, and across all stages GELU yields the lowest high/low ratios. This aligns with its smooth functional form, which naturally biases the network toward low-frequency representations. In contrast, ReLU and ReLU6 systematically amplify high-frequency components. In the early stages (Stage 0 and 1), ReLU exhibits the strongest high-frequency response, reflecting its non-smooth activation behavior and its tendency to preserve or enhance sharp transitions. In deeper layers (Stage 2 and 3), ReLU6 produces the highest high/low ratios, suggesting that its clipped nonlinearity becomes more influential as depth increases—potentially explaining why the ReLU6-based model obtains the best overall performance. These effects are stable across stages, blocks, and models, and they directly support our hypothesis: smooth activations such as GELU favor low-frequency features, whereas non-smooth activations (ReLU/ReLU6) amplify high-frequency content. This yields a clear falsifiable prediction—if a smooth activation were ever to systematically exceed ReLU/ReLU6 in high-frequency ratios under comparable settings, our hypothesis would be invalidated—which strengthens the explanatory robustness of the spectral analysis presented in Sec. 3.2.
A.11 Extra experiments on gating mechanisms.
In this section, we present additional experimental results demonstrating how different activation functions affect frequency learning. Firstly, we have conducted experiments with different settings of $r$ of Fig. 3 in the main paper. We have set $r$ to $\{4,7\}$ respectively. The results are shown in Fig., the performances are aligned with our analysis in Sec. 3 which indicates that the results shown in Fig. 3 are not accidental.
Specifically, we conduct experimearents using another smooth activation function, Swish (SiLU), and a non-smooth function, ReLU6.
As shown in Figures 14(a) and 14(b), ReLU6 performs similarly to GELU. Although ReLU6 may introduce more high-frequency components, it does not perform as well as ReLU for two main reasons: (1) ReLU6 caps the activation values. It can reduce the model’s sensitivity to high-frequency components where those components are often associated with higher activation values. (2) The use of low-resolution images can adversely affect the performance in classifying high-frequency components, as finer details are lost, making it harder for the model to learn these features. Figures 14(c) and 14(d) present comparisons between SiLU (Swish) and ReLU, as well as between SiLU and GELU, respectively. The Res18-Gate-SiLU performs better on lower-frequency components, specifically in the range $r\in(0,r_{1})$. This indicates that SiLU has a greater smoothing effect on the information, encouraging the model to learn more effectively from lower frequencies.
Moreover, we investigated the impact of different training strategies to understand their effects on model performance. Specifically, as plotted in the Fig. 15, we replaced the optimizer with Adam, setting its learning rate to 0.001. While the training curves showed noticeable variations compared to the baseline setup, the overall performance differences among the model variants remained consistent. This consistency indicates that the observed behaviors are robust to changes in optimization strategies.
We also conduct experiments on setting the redii to different sets. As shown in Figs. 16 and 17, we set radii to $0,4,8,12,+\infty]$ and $0,8,16,24,+\infty]$ respectively. When the frequency intervals are too small, the differences between the methods become less pronounced, especially in the lower-frequency components. When the frequency intervals are too large, all models struggle to classify higher-frequency components. Although differences between models become more pronounced in the lower-frequency components, such as between GELU and ReLU activations, to better understand the training dynamics, it is necessary to examine the differences in the high-frequency components as well. Therefore, we decide to display the results of setting radii to $0,6,12,24,+\infty]$ in the main body of the paper to have a better understand of the training dynamic of different variants. These results provide valuable insights into the functionality of the gating mechanisms. They suggest that the interaction between the element-wise product and the activation functions is a general phenomenon.
A.12 Downstream tasks.
We further provide the results of downstream tasks of Object Detection, Instance Segmentation and Semantic Segmentation. Firstly, we conducted experiments on GmNet-S3 on MSCOCO 2017 with the Mask RCNN
Table 15: Object detection & Instance segmentation& Semantic segmentation. The latency is tested on iPhone 14 by Core ML Tools.
[…]
(ms) APbox $^{box}_{50} AP$$^{box}_{75}$ APmask $^{mask}_{50} AP$$^{mask}_{75}$ mIoU
EfficientFormer-L3 12.4 41.4 63.9 44.7 38.1 61.0 40.4 43.5
RepViT-M1.5 6.9 41.6 63.2 45.3 38.9 60.5 41.5 43.6
GmNet-S3 5.2 42.2 63.4 46.7 40.1 61.2 42.9 44.6
[…]
A.13 The Usage of Large Language Models (LLMs)
We used GPT for polishing grammar and improving readability. All research ideas and analyses were conducted by the authors, who take full responsibility for the content.
(a) Comparison between Res18-Gate-GELU and Res18-Gate-ReLU6.
(b) Comparison between Res18-Gate-ReLU and Res18-ReLU6.
(c) Comparison between Res18-Gate-GELU and Res18-Gate-SiLU.
(d) Comparison between Res18-Gate-ReLU and Res18-Gate-SiLU.
Figure 14: Learning curves of Resnet18 and its variants for $100$ epoch. The radii are set to $\{0,6,12,18,+\infty\}$
(a) Comparison between Res18 and Res18-Ewp.
(b) Comparison between Res18-Gate-ReLU and Res18-Ewp.
(c) Comparison between Res18-Gate-ReLU and Res18-Gate-GELU.
Figure 15: Learning curves of Resnet18 and its variants for $100$ epochs with optimizer of Adam. The learning rate is set to $0.001$. Radii are set to $\{0,6,12,18,+\infty\}$
(a) Comparison between Res18 and Res18-Ewp.
(b) Comparison between Res18-Gate-ReLU and Res18-Ewp.
(c) Comparison between Res18-Gate-ReLU and Res18-Gate-GELU.
Figure 16: Learning curves of Resnet18 and its variants for $100$ epochs with optimizer of SGD. Radii are set to $\{0,4,8,12,+\infty\}$. When the frequency intervals are too small, the differences between the methods become less pronounced, especially in the lower-frequency components. However, it remains evident that the different variants have distinct effects on learning the various frequency components.
(a) Comparison between Res18 and Res18-Ewp.
(b) Comparison between Res18-Gate-ReLU and Res18-Ewp.
(c) Comparison between Res18-Gate-ReLU and Res18-Gate-GELU.
Figure 17: Learning curves of Resnet18 and its variants for $100$ epochs with optimizer of SGD. Radii are set to $\{0,8,16,24,+\infty\}$. When the frequency intervals are too large, all models struggle to classify higher-frequency components. Although differences between models become more pronounced in the lower-frequency components, such as between GELU and ReLU activations, to better understand the training dynamics, it is necessary to examine the differences in the high-frequency components as well.